OpenStudy (anonymous):

5-2/3x=1 please show work

OpenStudy (anonymous):

5 - 2/3x = 1 2/3x = 5 - 1 2/3x = 4 2 = 12x x = 2/12 x = 1/6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome!